# Fully Dynamic Euclidean $k$-Means

### v2

@@ 12 Augmented $k$-Means
$\Pr[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}]\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \Pr\bigg{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
$\leq\leavevmode\nobreak\ \Pr\bigg{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
$\leq\leavevmode\nobreak\ \frac{1}{\mathrm{poly}(\Delta^{d})}.$
@@
In this part of the paper, we provide our algorithm, its implementation, and its analysis. We start by explaining the framework of [BCF25a] in Section 13, with some major modifications that are necessary to ensure that we can implement it using our geometric data structures (see Part II). Next, in Section 14, we show how to implement this modified framework in Euclidean spaces with an update time of $\tilde{O}(n^{\epsilon})$. Finally, in Sections 15, 16, 17 and 18, we provide the analysis of the algorithm.
@@
Throughout the rest of the paper, we use notation $\tilde{O}$ to hide $\mathrm{poly}(d\epsilon^{-1}\log(n\Delta))$. $d$ is the dimension of the space. $n$ is the maximum size of the space, and $[0,\Delta]$ is the range of the coordinates of the points. Note that the aspect ratio of the space would be $\sqrt{d}\Delta$ since we only consider points in the grid $[\Delta]^{d}$.
@@
Throughout Sections 14, 15, 16, 17 and 18, we assume that the dimension of the input space $d$ is $O(\log n)$. Hence, the running time performance of all the data structures in Part II becomes $\tilde{O}(n^{\epsilon})$. The assumption $d=O(\log n)$ can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume that the input space is $\mathbb{R}^{m}$. We can initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$, and whenever we have an insertion of a point $x\in\mathbb{R}^{m}$, we apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$. Then, we consider feeding this new dynamic space into our algorithm. Note that according to this transformation, we will have an additive $\tilde{O}(m)$ in the update time of the algorithm, but we omit to write this down from now on since it only appears because of this preprocessing step, and the rest of the algorithm will not depend on that.
13 The Algorithmic Framework of [BCF25a]
@@
In this section, we explain how the algorithmic framework of [BCF25a] works, with some significant modifications that are necessary for our purpose. We start by defining the modified version of the main tool used in the algorithm in Section 13.1, which is called the robust solution. Then, we continue with the description of the algorithm in Section 13.2.
@@
Although the algorithm of [BCF25a] is primarily designed for $k$-median, it also works for $k$-means. The only consideration is that the analysis of [BCF25a] depends on the fact that the integrality gap of the LP-relaxation for $k$-median is constant. Since this fact is also correct for $k$-means (see Theorem 1.3 in [ANFSW20]), we can extend the analysis of [BCF25a] for $k$-means. The constant factors in the analysis might change, and here we provide a thorough analysis of the algorithm specifically for Euclidean $k$-means.
@@ 13.1 Robust Solution
Before we explain the algorithm of [BCF25a], we start with the necessary definitions for describing the algorithm. Since our goal is to achieve an update time of $\tilde{O}(n^{\epsilon})$, we slightly changed some of the structural definitions of [BCF25a] that enable us to work with approximate balls around points instead of exact balls.
@@
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
@@
$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}.$ (42)
@@ 13.2 Description of The Algorithm
In this section, we describe the algorithm framework of [BCF25a]. For simplicity. Later in Section 14, we provide our algorithm using this framework in Euclidean spaces that has a small update time of $\tilde{O}(n^{\epsilon})$. Note that since we are considering new Definition 13.2 and we will implement subroutines of the algorithm approximately, the parameters of the algorithm here change significantly. The constant values in different parts of the algorithm are extremely tied together. For simplicity, in this section, we use big $O$ notation to describe their algorithm, which makes it easier to read. We will provide the exact parameters in Section 14 and the proper proofs in Section 15.
@@ 14.1 Implementing Steps of an Epoch
[…]
@@
^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot\lambda^{51}}\right\rfloor.$
With some extra calculations, with the same argument as in [BCF25a], we can show that $\ell+1=\tilde{\Omega}(\widehat{\ell}+1)=\tilde{\Omega}(\ell^{\star}+1)$ and
@@ 14.2 Implementing Robustify
In our implementation, for each center $u\in S$ in our main solution, we always maintain an integer $t[u]$ that indicates the center $u$ is $t[u]$-robust, i.e. we want to maintain the variant that $t[u]$ is always greater than or equal to the smallest integer $t$ satisfying $\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}$ (see Condition (42)). Hence, to check whether a center $u$ satisfies Condition (42), it suffices to check whether $t[u]\geq t$ [^16] Note that there might be a center $u$ such that $t[u]<t$, but $u$ is still $t$-robust. More precisely, we maintain an approximation of $\mathrm{dist}(u,S-u)$ using Lemma 7.3 and find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$.
@@ 14.2.1 Identifying Non-Robust Centers
Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop. Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
@@
[…]
3:$S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$Foreach $u\in W_{1}\cup W_{2}$
[…]
5:$u\leftarrow$pop$(\mathcal{Y})$
[…]
@@ 14.2.2 Make-Robust
[…]
4:$x_{j-1}\leftarrow x_{j}$Else
[…]
6:$u_{0}\leftarrow x_{0}$
7:$S\leftarrow S-u+u_{0}$
8:Save $t[u_{0}]\leftarrow t$ together with $u_{0}$
@@ 15.1.2 Well-Separated Pairs
Lemma 15.6 ($k$-Means Version of Lemma 4.2 [BCF25a]).
Let $r\geq 0$ and $m\in[0,k]$. Consider any two sets of centers $S,V\subseteq X$ such that $|S|=k$ and $|V|=k+r$. If the number of well-separated pairs w.r.t. $(S,V)$ is $k-m$, then there exists a subset $\tilde{S}\subseteq S$ of size at most $k-\lfloor(m-r)/4\rfloor$ such that
@@ 15.1.3 Other Key Lemmas
Lemma 15.9 ($k$-Means Version of Double-Sided Stability (Lemma 4.1 [BCF25a])).
@@ 16.1 Approximation Ratio During the Epoch
We will show that throughout the epoch the solution is always a $\lambda^{53}$ approximation w.r.t. the current dataset. According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Step 1 of our algorithm (see Algorithm 6, Algorithm 6), we infer that
@@
Finally, since =^$\ell=\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor\leq\widehat{\ell}$, we have
@@ 16.2 Approximation Ratio At the End of the Epoch
Let $V\subseteq[\Delta]^{d}$ be an optimal $(k+\ell+1)$-median solution for the dataset $X^{(0)}$, i.e., $\left|V\right|=k+\ell+1$ and $\textsc{cost}(X^{(0)},V)=\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$. Let $m\in[0,k]$ be the unique integer such that there are $(k-m)$ well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},T\right)$. Let $\{(u_{1},v_{1}),(u_{2},v_{2}),\ldots,(u_{k-m},v_{k-m})\}\subseteq S_{\textnormal{{init}}}\times V$ denote the collection of $(k-m)$ well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},V\right)$. Define the set $\tilde{A}:=V\setminus\{v_{1},\ldots,v_{k-m}\}$. It is easy to verify that:
@@
$\leq p_{\epsilon}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})$
@@
We apply Lemma 15.6, by setting $r=\ell+1$, $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$. This implies the existence of a set $\tilde{S}\subseteq S_{\textnormal{{init}}}$ of at most $(k-b)$ centers, with $b=\lfloor(m-\ell-1)/4\rfloor$, such that
$\textsc{cost}(X^{(0)},\tilde{S})$$\leq$ $81\lambda^{40}\cdot\left(\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})+\textsc{cost}(X^{(0)},V)\right)$
@@
In the above derivation, the second inequality follows from the initial assumption at the beginning of the epoch (52), and the last inequality holds because $\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)\leq\mathrm{OPT}_{k}(X^{(0)})$ and $\lambda=(p_{\epsilon})^{2}$ is a large number (see Part III). Since $\mathrm{OPT}_{k-b}\left(X^{(0)}\right)\leq\textsc{cost}(X^{(0)},\tilde{S})$, we get
@@
Recall the way $\ell,\widehat{\ell}$ and $\ell^{\star}$ are defined in Step 1 of our algorithm and our implementation (Algorithm 6). From (64), it follows that $b\leq\ell^{\star}$. According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Algorithm 6, for $s_{i}=2s_{i-1}=2\widehat{\ell}$, from (Algorithm 6), we have
@@
Now, we have $b\geq\frac{m-\ell-1}{4}-1=\frac{m-\ell-5}{4}$ and ^$\ell\geq\frac{\widehat{\ell}}{36\cdot\lambda^{51}}-1$, we get
@@ 17 Recourse Analysis of Our Algorithm
The amortized recourse of the algorithm is at most $\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$, where $d$ is the dimension of the space, $n$ is the upper bound on the number of the points in the data-set, and $[0,\Delta]$ is the range of the coordinates of the points.

### v3

@@ 13 Augmented $k$-Means
$\Pr[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}]\penalty 10000\$$\leq\penalty 10000\ \Pr\bigg[|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \bigg|\penalty 10000\ \neg\mathcal{E}_{i}\bigg]$
$\leq\penalty 10000\ \Pr\bigg[\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \bigg|\penalty 10000\ \neg\mathcal{E}_{i}\bigg]$
$\leq\penalty 10000\ \frac{1}{\mathrm{poly}(\Delta^{d})}.$
@@
In this part of the paper, we provide our algorithm, its implementation, and its analysis. We start by explaining the framework of [BCF24] in Section 14, with some major modifications that are necessary to ensure that we can implement it using our geometric data structures (see Part II). Next, in Section 15, we show how to implement this modified framework in Euclidean spaces with an update time of $\tilde{O}(n^{\epsilon})$. Finally, in Sections 16, 17, 18 and 19, we provide the analysis of the algorithm.
@@
Throughout the rest of the paper, we use notation $\tilde{O}$ to hide $\mathrm{poly}(d\epsilon^{-1}\log(n\Delta))$. $d$ is the dimension of the space. $n$ is the maximum size of the space, and $0,\Delta]$ is the range of the coordinates of the points. Note that the aspect ratio of the space would be $\sqrt{d}\Delta$ since we only consider points in the grid $\Delta]^{d}$.
@@
Throughout Sections 15, 16, 17, 18 and 19, we assume that the dimension of the input space $d$ is $O(\log n)$. Hence, the running time performance of all the data structures in Part II becomes $\tilde{O}(n^{\epsilon})$. The assumption $d=O(\log n)$ can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume that the input space is $\mathbb{R}^{m}$. We can initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$, and whenever we have an insertion of a point $x\in\mathbb{R}^{m}$, we apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $\Delta]^{d}$. Then, we consider feeding this new dynamic space into our algorithm. Note that according to this transformation, we will have an additive $\tilde{O}(m)$ in the update time of the algorithm, but we omit to write this down from now on since it only appears because of this preprocessing step, and the rest of the algorithm will not depend on that.
14 The Algorithmic Framework of [BCF24]
@@
In this section, we explain how the algorithmic framework of [BCF24] works, with some significant modifications that are necessary for our purpose. We start by defining the modified version of the main tool used in the algorithm in Section 14.1, which is called the robust solution. Then, we continue with the description of the algorithm in Section 14.2.
@@
Although the algorithm of [BCF24] is primarily designed for $k$-median, it also works for $k$-means. The only consideration is that the analysis of [BCF24] depends on the fact that the integrality gap of the LP-relaxation for $k$-median is constant. Since this fact is also correct for $k$-means (see Theorem 1.3 in [integrality-gap-k-means]), we can extend the analysis of [BCF24] for $k$-means. The constant factors in the analysis might change, and here we provide a thorough analysis of the algorithm specifically for Euclidean $k$-means.
@@ 14.1 Robust Solution
Before we explain the algorithm of [BCF24], we start with the necessary definitions for describing the algorithm. Since our goal is to achieve an update time of $\tilde{O}(n^{\epsilon})$, we slightly changed some of the structural definitions of [BCF24] that enable us to work with approximate balls around points instead of exact balls.
@@
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
@@
[…]
@@ 14.2 Description of The Algorithm
In this section, we describe the algorithm framework of [BCF24]. For simplicity. Later in Section 15, we provide our algorithm using this framework in Euclidean spaces that has a small update time of $\tilde{O}(n^{\epsilon})$. Note that since we are considering new Definition 14.2 and we will implement subroutines of the algorithm approximately, the parameters of the algorithm here change significantly. The constant values in different parts of the algorithm are extremely tied together. For simplicity, in this section, we use big $O$ notation to describe their algorithm, which makes it easier to read. We will provide the exact parameters in Section 15 and the proper proofs in Section 16.
@@ 15.1 Implementing Steps of an Epoch
1: for $i=0$ to $\log_{2}k$ do Using restricted $k$-means (Lemma 8.5), in $\tilde{O}(n^{\epsilon}s_{i})$ time compute a subset $\widehat{S}_{i}\subseteq S_{\textnormal{{init}}}$ of $(k-s_{i})$ centers, that is a $p_{\epsilon}$-approximation to $\mathrm{OPT}_{k-s_{i}}^{S_{\textnormal{{init}}}}\left(X^{(0)}\right)$.
2:   if $\textsc{cost}(X^{(0)},\widehat{S}_{i})>(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})$ then
3:    return $\widehat{\ell}:=s_{i-1}$.
@@
$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot\lambda^{51}}\right\rfloor.$
With some extra calculations, with the same argument as in [BCF24], we can show that $\ell+1=\tilde{\Omega}(\widehat{\ell}+1)=\tilde{\Omega}(\ell^{\star}+1)$ and
@@ 15.2 Implementing Robustify
In our implementation, for each center $u\in S$ in our main solution, we always maintain an integer $t[u]$ that indicates the center $u$ is $t[u]$-robust, i.e. we want to maintain the variant that $t[u]$ is always greater than or equal to the smallest integer $t$ satisfying $\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}$ (see Condition (43)). Hence, to check whether a center $u$ satisfies Condition (43), it suffices to check whether $t[u]\geq t$. [^16] Note that there might be a center $u$ such that $t[u]<t$, but $u$ is still $t$-robust. More precisely, we maintain an approximation of $\mathrm{dist}(u,S-u)$ using Lemma 8.3 and find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$.
@@ 15.2.1 Identifying Non-Robust Centers
[…]
@@
1: $W_{1}\leftarrow W^{\prime}\setminus S_{\textnormal{{init}}}$ $\triangleright$ All new centers
2: for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ do
3:   for each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$ do
4:    Let $u\leftarrow\textsc{NN}(x,S[i])$
5:    if $\mathrm{dist}(u,x)\leq\lambda^{3i+1.5}$ then $W_{2}\leftarrow W_{2}+u$ $\triangleright$ Mark $u$ as a contaminated center
6: $S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$
7: for each $u\in W_{1}\cup W_{2}$ do
8:   $\texttt{MakeRobust}(u)$.
9: while $\mathcal{Y}\neq\emptyset$ do
10:   $u\leftarrowpop$$(\mathcal{Y})$
11:   $t\leftarrow$ Smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$
12:   if $t[u]\geq t$ then
13:    go back to 9
14:   else
15:    $\texttt{MakeRobust}(u)$.
16: return $S_{\textnormal{{final}}}\leftarrow S$ .
@@ 15.2.2 Make-Robust
2: $x_{t}\leftarrow u$
3: for $j=t$ down to $1$ do
4:   Invoke Lemma 8.4 for $x=x_{j}$ and $r=\lambda^{3j}$ to get $c^{*}\in[\Delta]^{d}$, $\widehat{\textsc{cost}}_{x_{j}}$, $\widehat{\textsc{cost}}_{c^{*}}$ and $b_{x_{j}}$
5:   if $\left(\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq\lambda^{3j-1}/p_{\epsilon}\right)$ or $\left(\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{*}}\right)$ then
6:    $x_{j-1}\leftarrow x_{j}$
7:   else
8:    $x_{j-1}\leftarrow c^{*}$
9: $u_{0}\leftarrow x_{0}$
10: $S\leftarrow S-u+u_{0}$
11: Save $t[u_{0}]\leftarrow t$ together with $u_{0}$
@@ 16.1.2 Well-Separated Pairs
Lemma 16.6 ($k$-Means Version of Lemma 4.2 [BCF24]).
Let $r\geq 0$ and $m\in[0,k]$. Consider any two sets of centers $S,V\subseteq X$ such that $|S|=k$ and $|V|=k+r$. If the number of well-separated pairs w.r.t. $(S,V)$ is kmk-m, then there exists a subset $\tilde{S}\subseteq S$ of size at most $k-\lfloor(m-r)/4\rfloor$ such that
@@ 16.1.3 Other Key Lemmas
Lemma 16.9 ($k$-Means Version of Double-Sided Stability (Lemma 4.1 [BCF24])).
@@ 17.1 Approximation Ratio During the Epoch
We will show that throughout the epoch the solution is always a $\lambda^{53}$ approximation w.r.t. the current dataset. According to the definition of $\widehat{\ell}=s_{i-1}$ in Step 1 of our algorithm (see Algorithm 6, 2), we infer that
@@
Finally, since $\ell=\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor\leq\widehat{\ell}$, we have
@@ 17.2 Approximation Ratio At the End of the Epoch
Let $V\subseteq[\Delta]^{d}$ be an optimal $(k+\ell+1)$-median solution for the dataset $X^{(0)}$, i.e., $\left|V\right|=k+\ell+1$ and $\textsc{cost}(X^{(0)},V)=\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$. Let $m\in[0,k]$ be the unique integer such that there are (km)(k-m) well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},T\right)$. Let $\{(u_{1},v_{1}),(u_{2},v_{2}),\ldots,(u_{k-m},v_{k-m})\}\subseteq S_{\textnormal{{init}}}\times V$ denote the collection of (km)(k-m) well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},V\right)$. Define the set $\tilde{A}:=V\setminus\{v_{1},\ldots,v_{k-m}\}$. It is easy to verify that:
@@
[…]
@@
We apply Lemma 16.6, by setting $r=\ell+1$, $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$. This implies the existence of a set $\tilde{S}\subseteq S_{\textnormal{{init}}}$ of at most (kb)(k-b) centers, with $b=\lfloor(m-\ell-1)/4\rfloor$, such that
[…]
@@
[…]
@@
Recall the way $\ell,\widehat{\ell}$ and $\ell^{\star}$ are defined in Step 1 of our algorithm and our implementation (Algorithm 6). From (65), it follows that $b\leq\ell^{\star}$. According to the definition of $\widehat{\ell}=s_{i-1}$ in Algorithm 6, for $s_{i}=2s_{i-1}=2\widehat{\ell}$, from (2), we have
@@
Now, we have $b\geq\frac{m-\ell-1}{4}-1=\frac{m-\ell-5}{4}$ and $\ell\geq\frac{\widehat{\ell}}{36\cdot\lambda^{51}}-1$, we get
@@ 18 Recourse Analysis of Our Algorithm
The amortized recourse of the algorithm is at most $\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$, where $d$ is the dimension of the space, $n$ is the upper bound on the number of the points in the data-set, and $0,\Delta]$ is the range of the coordinates of the points.